# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ 5.1 Finding Influential Examples
Recall that we cannot access the sensitive attribute in datasets $P$ and $U$ due to privacy concerns. The absence of sensitive attributes makes it impossible to evaluate current fairness violations and ascertain the fairness of the classifier, when we lack any knowledge about the test dataset $Q$. To overcome this limitation, we first introduce an auxiliary hold-out validation set $Q_{v}:=\{(x_{n}^{\circ},y_{n}^{\circ},a_{n}^{\circ})\}_{n=1}^{N_{v}}$, drawing from the distribution $\mathcal{Q}$ (Kirichenko et al., 2022; Liu et al., 2021b). Please note that our access is restricted solely to the sensitive attributes presented in this validation set. Despite its small size, this validation dataset offers valuable insights that aid in achieving our ultimate goal.
Inspired by recent work (Diesendruck et al., 2020; Choi et al., 2020; Li and Vasconcelos, 2019; Pruthi et al., 2020), to address the lack of sensitive attributes, the high-level idea is that we can evaluate the impact of each example on fairness by comparing the gradient originating from a single data example with the gradient derived from the entire validation set. This comparison helps in quantifying the potential advantage of including this specific example in improving fairness. Intuitively, if the gradient obtained from a single data example has a similar direction to the gradient from the validation set, it indicates that incorporating this example contributes to enhancing the model’s fairness. Similarly, the negative impact of a single data example on prediction can also be avoided by using its gradient.
In this regard, an ideal new example can help reduce the fairness disparity while keeping accuracy non-decreasing. Here, we consider the loss for the prediction component and fairness component, i.e., $\ell(\mathbf{w},z)$ and $\phi(\mathbf{w},z)$. Without loss of generality, we assume that both the loss function $\ell(\cdot)$ and fairness loss $\phi(\cdot)$ are differentiable with respect to $\mathbf{w}$ in the optimization process. Suppose that the model is updated following the gradient decent, the change of model parameters by counterfactually optimized on a new instance $z^{\prime}$ is
@@
where $\eta$ refers to the learning rate. Following this, we compute the influence of the prediction and fairness components, respectively.
Influence of Prediction Component. The influence for the prediction component on the validation data $z_{n}^{\circ}=(x_{n}^{\circ},y_{n}^{\circ},a_{n}^{\circ})$, when model parameters is updated from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$ by adding a new example $z^{\prime}$, is given by
@@
The prediction influence of new example $z^{\prime}$ on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=$ $\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx$ $-\eta\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\sum_{n=1}^{N_{v}}\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$ (6)
Intuitively, the more negative $\textsf{Infl}_{\textsf{acc}}(z^{\prime})$ is, the more positive on the model prediction (performance) that example $z^{\prime}$ can provide.
Influence of Fairness Component. Recall that $\phi$ denotes the fairness loss. The fairness influence on validation data $z_{n}^{\circ}$ when model parameters change from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$ by adding $z^{\prime}$ is denoted by
@@
For simplicity, we write $\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ};\mathbf{w}_{t},\mathbf{w}_{t+1})$ as $\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})$. Then, similarly, we have
@@
The fairness influence of new example $z^{\prime}$ on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{fair}}(z^{\prime}):=$ $\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx$ $-\eta\sum_{n\in[N_{v}]}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$ (8)
Similar to the prediction component, the greater the negativity of $\textsf{Infl}_{\textsf{fair}}(z^{\prime})$ is, the greater the positive impact that the example $z^{\prime}$ has on fairness.
It is important to note that neither the influence of prediction nor fairness components require the annotations for the sensitive attributes of the sample $z^{\prime}$, which are unavailable in the dataset $U$. In contrast, calculating the first-order gradient of fairness loss $\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})$ only relies on the sensitive attributes of validation data $z^{\circ}$, which we assume to be available to the model learner.
We also validate how accurate the first-order estimation of the influence is in comparison to the real influence (Koh and Liang, 2017), and find that the estimated influences for most of examples are very close to their actual influence values. We refer the readers to Appendix D.3 for more details.
@@ 5.2 Algorithm: Fair Influential Sampling (FIS)
Drawing on Lemma 5.1 and Lemma 5.2, we can efficiently pinpoint those examples with the most negative fairness influence and negative prediction influence. This sampling method aids in reducing fairness disparities while sidestepping adverse impacts on accuracy.
Label Annotation. Before diving into presenting our sampling algorithm, it is necessary to address the problem of not having access to the true labels of new solicited examples. Lacking the label information for new examples poses a significant challenge in determining the corresponding influence on predictions and fairness, a fact that is substantiated by Lemma 5.1 and Lemma 5.2. In practice, one can always recruit human annotators to get the ground-truth labels for those unlabeled examples. However, we assume that it is still cost-ineffective.
To tackle this problem, one common approach is utilizing a model that has been effectively trained on dataset $P$ to produce proxy labels, which serve to approximate the calculation of influences for examples from a substantial unlabeled dataset $U$. It’s important to note that these proxy labels are exclusively used during the sampling phase. Here, we propose to annotate the proxy labels with the model trained on the labeled set $P$. In particular, we introduce a strategy that employs lowest-influence labels for the annotation of label $y^{\prime}$ given $x^{\prime}$:
$\hat{y}=\operatorname*{arg\,min}_{k\in[K]}~{}|\textsf{Infl}_{\textsf{acc}}(x^{\prime})|,$ (9)
which corresponds to annotating the most confident label class. After sampling, we still need to acquire the true labels of the selected data examples for subsequent training. We are aware that the true labels may not always be consistent with the elicited labels even with human annotations (Zhu et al., 2021, 2024; Cheng et al., 2020; Wei et al., 2022). The gap between human annotations and true labels is left for future works.
@@
1:  Input: train set $P$, unlabeled set $U$, validation set $Q_{v}$, $t=0$, number of new examples in each round $r$, number of rounds $T$, tolerance $\epsilon$# Training phrase: warm start ##
2:  Train classifier $f$ solely on $P$ by minimizing the empirical risk $R_{\ell}$. Obtain model parameters $\mathbf{w}_{1}$ and validation accuracy (on $Q_{v}$) $\text{VAL}_{0}$. ## Iterative sampling phrase ##
@@
4:     Guess label $\hat{y}$ for new examples using Eq. (9).
5:     Calculate the influence of prediction and fairness component by Eq. (6) and Eq. (8), respectively.
6:     Get the scores of fair examples:
$S_{\text{o}}=\{\textsf{Infl}_{\textsf{fair}}(z^{\prime})~{}|~{}\textsf{Infl}_{\textsf{acc}}(z^{\prime})\leq 0,\textsf{Infl}_{\textsf{fair}}(z^{\prime})\leq 0\}$
7:     $S_{t}=\{\}$
8:     while $|S_{t}|<r$ do
9:        Find top-$(r-|S_{t}|)$ annotated examples (lowest fairness influence):
$\{z^{\prime}_{n}\}\leftarrow\textsf{Top-}r(S_{\text{o}})$
@@
11:     end while## Training phrase: incorporate new examples ##
12:     Train model $f$ on $P\cup S_{1}\cdots\cup S_{t}$ with true inquired labels by minimizing $R_{\ell}$. Obtain model parameters $\mathbf{w}_{t+1}$ and validation accuracy (on $Q_{v}$) $\text{VAL}_{t}-\epsilon$.
13:  end forOutput: models $\{\mathbf{w}_{t}~{}|~{}\text{VAL}_{t}>\text{VAL}_{0}\}$
Proposed algorithm. Incorporating sampling techniques, we develop a Fair Influential Sampling (FIS) method to select a limited number of influential examples from dataset $U$. The full training algorithm for fair influential sampling is summarized in Algorithm 1. Note that the tolerance $\epsilon$ is applied to monitor the performance drop in validation accuracy. In Line 2, we initiate the process by training a classifier $f$ solely on dataset $P$, that is, performing a warm start. Subsequently, T-round sampling iterations are applied to amend dataset $P$, aiming to reduce fairness disparities. Following the iterative fashion, FIS guesses labels using the proposed strategy I in Line 4. In particular, in Lines 5–6, we calculate prediction’s and fairness’s influences for examples using Lemma 5.1 and Lemma 5.2, respectively. Then, the true labels of the top-$r$ samples with the most negative impact on fairness will be inquired for further training (Lines 8-11). In Line 12, new examples will be incorporated with the training dataset $P$, followed by the continued training of the fair classifier using datapoints sampled from the updated set $P\cup S_{1}\cdots\cup S_{t}$. To amplify the influence of these new datapoints, they are assigned a greater weight compared to the datapoints in the original training dataset $P$.
Note that although we propose a specific strategy for guessing labels, our algorithm is a flexible framework that is compatible with any other labeling method. In particular, we save the model parameters at each round as checkpoints $\mathbf{w}_{t}$. Ultimately, to avoid potential accuracy drops incurred by excessively large random perturbations, we exclusively choose and offer models for output whose validation accuracy exceeds the initial validation accuracy $\text{VAL}_{0}$.

### v3

@@ 4.1 Finding influential examples
To avoid using training sensitive attributes, our primary idea is to find newly acquired data that assists in creating a “fairer” dataset, which allows for training a fair and accurate model via standard ERM. Initially, we explore whether newly acquired data enhances fairness by examining the training process, where the model is typically updated using gradient descent. The change of model parameters by performing one step gradient descent on newly acquired data $z^{\prime}$ is
@@
where $\eta$ refers to the learning rate and $\ell(\cdot)$ is the training loss function. It should be noted that before we solicit the true labels of samples $z^{\prime}$, we first use proxy labels. In the following subsection 4.2, we will present a strategy for proxy labels. Training on $z^{\prime}$ affects the model’s prediction on validation data $z_{n}^{\circ}$ regarding both accuracy and fairness. If the updated model $\mathbf{w}_{t+1}$ outperforms the previous one $\mathbf{w}_{t}$ evaluated on the validation dataset in terms of fairness and accuracy, this acquired data $z^{\prime}$ helps to reduce the fairness disparity without worsening accuracy.
To separately measure the accuracy and fairness performance of the updated model on the validation set, we introduce two types of loss functions: fairness loss $\underline{\phi(\mathbf{w},z_{n}^{\circ})}$ and accuracy loss $\underline{\ell(\mathbf{w},z_{n}^{\circ})}$, where validation data $z_{n}^{\circ}=(x_{n}^{\circ},y_{n}^{\circ},s_{n}^{\circ})$. Note that these loss functions are developed for sampling, not for training. Besides, training loss function $\ell(\cdot)$ can be reused as the accuracy loss function due to the same update target. One can identify training loss $\ell(\cdot,z^{\prime})$ and accuracy loss $\ell(\cdot,z_{n}^{\circ})$ based on the input data used. Without loss of generality, we assume that $\ell(\cdot)$ and $\phi(\cdot)$ are differentiable w.r.t. $\mathbf{w}$. Here, we do not restrict the generality of the fairness loss function; it can be any smoothed version of fairness metrics such as DP or EOd. Following this, we develop the influence of the accuracy and fairness components for finding the samples, respectively.
Influence of accuracy component
When model parameters are updated from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$ by adding a new example $z^{\prime}$, the influence of model’s accuracy on one validation example $z_{n}^{\circ}$ is:
@@
The accuracy influence of new example $z^{\prime}$ on the validation dataset $Q_{v}$ is:
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=\sum\nolimits_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx-\eta\sum\nolimits_{n\in|Q_{v}|}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$ (2)
Intuitively, the more negative $\textsf{Infl}_{\textsf{acc}}(z^{\prime})$ is, the more positive the model accuracy (performance) that example $z^{\prime}$ can provide.
Influence of fairness component
When model parameters are updated from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$ by adding a new example $z^{\prime}$, the influence of model’s fainess on one validation example $z_{n}^{\circ}$ is:
@@
For simplicity, we write $\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ};\mathbf{w}_{t},\mathbf{w}_{t+1})$ as $\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})$. Then, similarly, we have:
@@
The fairness influence of new example $z^{\prime}$ on the validation dataset $Q_{v}$ is:
$\textsf{Infl}_{\textsf{fair}}(z^{\prime}):=\sum\nolimits_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx-\eta\sum\nolimits_{n\in|Q_{v}|}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$ (4)
Similar to the accuracy component, the greater the negativity of $\textsf{Infl}_{\textsf{fair}}(z^{\prime})$ is, the greater the positive impact that the example $z^{\prime}$ has on fairness.
Intuitions
These two components evaluate the accuracy and fairness impact of each example by comparing the gradient originating from a single data sample with the gradient derived from the entire validation set, respectively. This comparison helps quantify the potential advantage of including this specific example in training. For instance, if the gradient obtained from one example has a similar direction to the gradient from the validation set, it indicates that incorporating this example contributes to enhancing the model’s fairness or accuracy.
Training sensitive attributes are not disclosed
One can easily check that neither the influence of accuracy nor fairness components require the sensitive attributes of any example $z^{\prime}$, as the example $z^{\prime}$ only appears in the first-order gradient of the accuracy loss $\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime})$. In the fairness component, calculating the $\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})$ only relies on validation example $z^{\circ}_{n}$’s sensitive attributes. Here, we also validate how accurate the first-order estimation of the influence is in comparison to the real influence [39], and find that the estimated influences for most of the examples are very close to their actual influence values. We refer the readers to Appendix C.1 for more details.
Even without disclosing training sensitive attributes, the correlations between non-demographic features and demographic information may still lead to privacy leakage issues [82]. To address this potential privacy concern, we provide further discussions and theoretical analysis using differential privacy in Appendix F.
@@ 4.2 Algorithm: fair influential sampling (FIS)
Following Lemma 4.1 and Lemma 4.2, we can efficiently select those examples with the most negative fairness influence and negative accuracy influence. This sampling method aids in reducing fairness disparities without worsening model accuracy.
Labeling
Before presenting our sampling algorithm, it is necessary to address the problem of not accessing the true labels of new solicited examples. Lacking the label information for new examples poses a challenge in determining the corresponding influence on accuracy and fairness, a fact that is substantiated by Lemma 4.1 and Lemma 4.2. Intuitively, one can always recruit human annotators to get the ground-truth labels for those unlabeled examples. However, it is impractical due to the limited labeling budgets. To tackle this problem, another common approach is utilizing a model that has been effectively trained on dataset $P$ to produce proxy labels, which approximate the calculation of influences for examples from a substantial unlabeled dataset $U$. It’s important to note that these proxy labels are exclusively used during the sampling phase. To maintain good model performance, we still need to inquire about the true labels of the selected data examples for subsequent training. Here, we propose to annotate the proxy labels with the model trained on the labeled set $P$. In particular, we introduce a strategy that employs lowest-influence labels for annotating label $\hat{y}^{\prime}$ given $x^{\prime}$:
$\hat{y}^{\prime}=\operatorname*{arg\,min}_{k\in\{1,\cdots,K\}}~{}|\textsf{Infl}_{\textsf{acc}}(x^{\prime},k)|,$ (5)
Here, we denote $\hat{z}^{\prime}:=(x^{\prime},\hat{y}^{\prime})$ for the proxy labels.
@@
1:  Input: training set $P$, unlabeled set $U$, validation set $Q_{v}$, new acquired set $S_{t}=\{\},\forall t\in[T]$ rounds, number of new selected examples in each round $r$, tolerance $\epsilon$.
2:  Warmup: Train a classifier $f$ solely on $P$ by minimizing the empirical risk $R_{\ell}$. Obtain model parameters $\mathbf{w}_{1}$ and validation accuracy (on $Q_{v}$) $\text{VAL}_{0}$.
@@
4:     Guess proxy label $\hat{y}^{\prime}$ for new examples $\hat{z}^{\prime}$ using Eq. (5).
5:     Compute the influence of accuracy and fairness component using Eq. (2) and Eq. (4):
6:     $\qquad\qquad\qquad S_{\text{original}}=\{\textsf{Infl}_{\textsf{fair}}(\hat{z}^{\prime})~{}|~{}\textsf{Infl}_{\textsf{acc}}(\hat{z}^{\prime})\leq 0,\textsf{Infl}_{\textsf{fair}}(\hat{z}^{\prime})\leq 0,\hat{z}^{\prime}\in U\}$
7:     while $|S_{t}|<r$ do
8:        Find top-$(r-|S_{t}|)$ annotated examples $\hat{z}^{\prime}_{n}$ based on the lowest fairness influence and then inquire about true labels $y^{\prime}$:
9:        $\qquad\qquad\qquad\qquad\{z^{\prime}_{n}\}\stackrel{{\scriptstyle\text{inquiring}}}{{\longleftarrow}}\{\hat{z}^{\prime}_{n}\}\leftarrow\textsf{Top-$(r-|S_{t}|)$}(S_{\text{original}})$
@@
11:        $U\leftarrow U\cap S_{t};\quad S_{\text{original}}\leftarrow S_{\text{original}}\cap S_{t}$
12:     end while
13:     Continue to train the model $f$ on the set $S_{t}$ via standard ERM. Obtain the updated model parameters $\mathbf{w}_{t+1}$. If the model’s validation accuracy (on $Q_{v}$) $\text{VAL}_{t}$ does not meet the desired threshold $\text{VAL}_{0}$, reject the updated model.
14:  end for
15:  Output: models $\{\mathbf{w}_{t}~{}|~{}\text{VAL}_{t}>\text{VAL}_{0}-\epsilon\}$
Proposed algorithm
The full procedure is outlined in Algorithm 1. Note that the tolerance $\epsilon$ is applied to monitor the performance drop in validation accuracy. In Line 2, we initiate the process by training a classifier $f$ solely on dataset $P$, that is, performing a warm start. Subsequently, $T$-round sampling iterations are applied to acquire more examples to dataset $P$. Following the iterative fashion, FIS guesses labels using Eq. (5) in Line 4. Then, we calculate the scores for proxy examples based on the accuracy and fairness influence using Eq. (2) and Eq. (4) respectively. In Lines 6-12, we would opt for $r$ samples based on the influence scores, and inquire about the true labels of these examples. However, due to the gap between proxy labels $\hat{y}^{\prime}$ and true label $y$, the accuracy and fairness influences of the top-$r$ samples based on inquired true labels $z^{\prime}$ may not necessarily satisfy the same conditions ($\textsf{Infl}_{\textsf{acc}}(\hat{z}^{\prime})\leq 0,\textsf{Infl}_{\textsf{fair}}(\hat{z}^{\prime})\leq 0$). Therefore, we use a while loop to iteratively select the top-$(r-|S_{t}|)$ examples for labeling until we obtain $r$ samples whose fairness influences based on true labels meet the conditions. Subsequently, in Line 11, we update sets $U$ and $S_{\text{original}}$ to prevent duplicate sampling. In Line 13, we would continue training using new examples with true inquired labels from set $S_{t}$. We save the model parameters at each round as checkpoints $\mathbf{w}_{t}$. To avoid potential accuracy drops incurred by excessively large random perturbations, we exclusively choose and offer models for output whose validation accuracy exceeds the initial validation accuracy $\text{VAL}_{0}$. Although we propose a specific strategy for guessing labels, our algorithm is flexible and compatible with other labeling methods. A comparative analysis of computational costs is detailed in Appendix C.2.
5 How more data improves fairness without harm?
In general, acquiring new data to supplement the original training dataset would potentially raise the distribution shift problem, affecting both accuracy and fairness. In this section, from a distribution shift perspective view, we first present a generalization error bound (accuracy side, Theorem 5.1) and risk disparity bound (fairness side, Theorem 5.2). The theoretical results jointly provide a high-level key insight that controlling the negative impact of distribution shift on generalization error, which refers to the model accuracy, could allow for improving fairness without harm. This theoretical insight validates the possibility of our sampling approach.
Without loss of generality, we discretize the whole distribution space and suppose that the train/test distributions are both drawn from a series of component distributions $\{\pi_{1},\cdots,\pi_{I}\}$ [28]. Then, the empirical risk $\mathcal{R}_{P}(\mathbf{w})$ calculated over a training set $P$ can be reformulated by splitting samples based on the component distributions:
$\mathcal{R}_{P}(\mathbf{w}):=\mathbb{E}_{z\in P}[\ell(\mathbf{w},z)]=\sum\nolimits_{i=1}^{I}p^{(P)}(\pi=i)\cdot\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w},z)].$
where $p^{(P)}(\pi=i)$ represents the frequencies of examples in $P$ drawn from component distribution $\pi_{i}$. Then, we can define the measure of probability distance between two sets or distributions as $\text{dist}(\mathcal{A},\mathcal{B}):=\sum_{i=1}^{I}|p^{(\mathcal{A})}(\pi=i)-p^{(\mathcal{B})}(\pi=i)|$. To reflect the implicit unfairness in the models, we introduce two basic assumptions in convergence analysis [43].
Assumption 5.1 ($L$-Lipschitz Continuous).
There exists a constant $L>0$, for any $\mathbf{v},\mathbf{w}\in\mathbb{R}^{d}$, $\mathcal{R}_{P}(\mathbf{v})\leq\mathcal{R}_{P}(\mathbf{w})+\langle\nabla\mathcal{R}_{P}(\mathbf{w}),\mathbf{v}-\mathbf{w}\rangle+\frac{L}{2}\lVert\mathbf{v}-\mathbf{w}\rVert^{2}_{2}$.
Assumption 5.2 (Bounded Gradient on Random Sample).
The stochastic gradients on any sample $z$ are uniformly bounded, i.e., $\mathbb{E}[||\nabla\mathcal{R}_{P}(\mathbf{w}_{t},z)||^{2}]\leq G^{2}$, and training epoch $t\in[1,\cdots,T]$.
Analogous to Assumption 5.2, we further make a mild assumption to bound the loss over the component distributions $\pi_{i}$ according to the corresponding model, that is, $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$, where $G_{P}$ is a bounding constant. For completeness, we first analyze the upper bound of generalization error, specifically from the standpoint of distribution shifts. Omitted proof can be found in Appendix D.
Theorem 5.1 (Generalization error bound).
Let $\text{dist}(\mathcal{P},\mathcal{Q})$, $G_{P}$ be defined therein. With probability at least $1-\delta$ with $\delta\in(0,1)$, the generalization error bound of the model trained on dataset $P$ is
$\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})\leq\underbrace{G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})}_{\text{ distribution shift}}+\sqrt{\frac{\log(4/\delta)}{2|P|}}+\mathcal{R}_{P}(\mathbf{w}^{P}).$
Note that the generalization error bound is predominantly impacted by the shift in distribution, especially when we consider an overfitting model, i.e., the empirical risk $\mathcal{R}_{P}(\mathbf{w}^{P})\to 0$.
Theorem 5.2 (Upper bound of risk disparity).
Suppose $\mathcal{R}_{\mathcal{Q}}(\cdot)$ follows Assumption 5.1. Let $\text{dist}(\mathcal{P},\mathcal{Q})$, $G_{P}$, $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})$ and $\text{dist}(P_{k},P)$ be defined therein. The initial learning rate $\eta_{0}$ satisfies $\eta_{0}^{2}<\frac{1}{\sqrt{2}TL}$, where $T$ is the number of training epochs. Given model $\mathbf{w}^{P}$ and $\mathbf{w}^{k}$ trained exclusively on group $k$’s data $P_{k}$, with probability at least $1-\delta$ with $\delta\in(0,1)$, then the upper bound of risk disparity is
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})\leq\underbrace{G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})}_{\text{ distribution shift}}+\underbrace{4L^{2}G^{2}\cdot\text{dist}(P_{k},P)^{2}}_{\text{group gap}}+\Upsilon$ (6)
where $\Upsilon=\sqrt{\frac{\log(4/\delta)}{2|P|}}+\sqrt{\frac{\log(4/\delta)}{2|P_{k}|}}+\varpi+\varpi_{k}.$ Note that $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k}$, $\varpi=\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w}^{\mathcal{Q}})$ and $\varpi_{k}=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$. $\varpi$ and $\varpi_{k}$ can be regarded as constants because $\mathcal{R}_{P}(\mathbf{w}^{P})$ and $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ correspond to the empirical risks, $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ represent the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$ and $\mathcal{Q}_{k}$, respectively.
Interpretations of Theorem 5.2
Eq. (6) illustrates several aspects that induce unfairness. (1) Group biased data. For group-level fairness, the more balanced the data is, the smaller the risk disparity would be; (2) distribution shift. For source/target distribution, the closer the distributions are, the smaller the performance gap would be; (3) Data size. For training data size, a larger data size (potentially eliminating data bias across groups) would lead to a smaller performance gap.
Main observation
Theorem 5.1 underscores how the generalization error is impacted by distribution shifts. Theorem 5.2 implies that the risk disparity is essentially influenced by the distribution shift and the inherent group gap term. In practice, approaches that mitigate the group gap, such as imposing fairness regularizers, acquiring new data, or reweight the training data samples [37], will inevitably incur additional distribution shifts between the training and test data. The incurred distribution shift further leads to a performance drop due to the generalization error in Theorem 5.1. Nonetheless, one theoretical insight is that if one can control the negative impacts of potential distribution shifts through generalization error while implementing fairness-enhancing strategies, it is possible to achieve the goal of improving fairness without causing harm. This high-level insight supports the effectiveness of our proposed sampling approach, in which we acquire new data to reduce the group gap through fairness components while preventing the potential adverse impacts of distribution shifts using the accuracy influence component.